What is the t value of the distribution with a= 1% and n=8 under the one tailed distribution

1
Expert's answer
2022-06-06T05:42:39-0400

Based on the information provided, the significance level is "\alpha = 0.01," "df=n-1=7" 

The critical value for a one-tailed left-tailed test is "t_c =-2.99793."

The critical value for a one-tailed right-tailed test is "t_c =2.99793."